# GPU Memory and Utilization Estimation for Training-Aware Resource Management: Opportunities and Limitations

### v1

@@ Abstract.
Collocating deep learning training tasks improves GPU utilization but causes drastic slowdowns due to resource contention and risks Out-of-Memory (OOM) failures. Accurate memory estimation is essential for robust collocation, while GPU utilization — a key proxy for resource contention — enables interference-aware scheduling to reduce slowdowns and improve throughput. Existing GPU memory estimators span three paradigms — analytical models, CPU-side libraries, and ML-based estimators — each with distinct limitations: dependence on detailed model specifications, intrusive integration, poor generalization, and varying latency overhead. GPU heterogeneity further complicates estimation, as identical tasks can exhibit markedly different memory footprints across hardware generations. GPU utilization remains comparatively understudied, further complicated by the non-additive nature of utilization metrics and hardware sensitivity.
We conduct a systematic analysis of representative estimators from each paradigm — Horus (Yeung et al., 2022), PyTorch FakeTensor (Contributors, 2025), and our lightweight ML-based estimator — evaluating accuracy, generalizability, and practical overhead. We construct a synthetic dataset spanning MLPs, CNNs, and Transformers with controlled architectural variations, and train MLP- and Transformer-based estimators for memory prediction. We further experiment with utilization estimation on the same dataset. Our evaluation reveals key tradeoffs and validates estimators against real-world unseen models. Significant challenges remain: analytical models are hardware-dependent, CPU-side libraries impose intrusive integration costs, and ML-based estimators struggle with cross-architecture generalization. We release all datasets, tools, and artifacts to support further research.
@@ 1. Introduction
Accurate GPU memory estimation is critical for robust resource management in deep learning workloads, enabling appropriate GPU selection and safe task collocation without exceeding memory limits (Robroek et al., 2024). Estimating related metrics such as GPU utilization, as a proxy for interference, and memory bandwidth can further optimize collocation decisions and reduce performance degradations because of resource contention. However, estimation introduces latency into the critical path of resource manager decision-making, necessitating fast estimation methods, offloading to machine learning frameworks during development, or pre-computation before tasks reach the mapper component.
@@
Deep learning training tasks are compute-intensive and demand substantial GPU memory based on model architecture, model size, input data size, and batch size, making efficient memory management essential. Maximizing GPU utilization is also necessary for energy efficiency (Jeon et al., 2019; Gao et al., 2024; Weng et al., 2022). Modern frameworks such as PyTorch and TensorFlow (Paszke et al., 2019; Abadi et al., 2016) employ optimization techniques like activation reuse and dynamic allocation, introducing unpredictability into memory usage patterns of training tasks. Consequently, simple estimation formulas based solely on numbers of parameters, gradients and activations often fail to produce accurate predictions. This complexity is further compounded by the fact that identical training tasks can exhibit markedly different memory footprints across different GPU hardware models (Albahar et al., 2022), owing to differences in software stacks, and hardware architecture as discussed in Section 2. Beyond memory, GPU utilization estimation faces distinct challenges: hardware heterogeneity causes the same task to exhibit different utilization across GPU architectures, and utilization metrics are non-additive, complicating collocation decisions.
@@
Existing GPU memory estimation approaches — spanning analytical methods (Yeung et al., 2022; Gao et al., 2020), library-based CPU-side tools (Contributors, 2025; Shi et al., 2025), and ML-based estimators (Gao et al., 2023) — universally suffer from over- or underestimation, dependence on detailed model specifications or computation graph, poor generalizability, and latency overhead on the critical scheduling/mapping path. GPU utilization estimation remains comparatively understudied, despite its controversial interpretation as a proxy for interference-aware collocation. Reactive and/or profiling-based collocation systems such as Gandiva (Xiao et al., 2018), Gavel (Narayanan et al., 2020), and Orion (Strati et al., 2024) each address interference management under different assumptions, yet share fundamental limitations in intrusiveness, scheduling granularity, or profiling overhead. These challenges can further get compounded if workloads show a change in their behavior, where memory consumption and GPU utilization can shift substantially over a job’s execution lifetime.
In this paper, we analyze existing GPU memory and utilization estimators, expose their practical limitations, and also release GPUMemNet and GPUUtilNet [^1] All artifacts, datasets, and reproducibility scripts are publicly available at https://github.com/itu-rad/GPUMemNet and https://github.com/itu-rad/GPUUtilNet., deep learning-based estimators for GPU memory consumption and utilization metrics respectively, achieving up to 97% estimation accuracy for MLPs and 82–87% for CNN and Transformer models, with limited generalization to real-world architectures such as ResNet, EfficientNet, and BERT.
Our contributions are as follows:
@@
  • We organize existing GPU memory estimation methods into three categories — analytical, library-based, and ML-based — and provide a comprehensive qualitative and quantitative comparison of representative techniques from each class: Horus (Yeung et al., 2022), PyTorch FakeTensor (Contributors, 2025), and our own GPUMemNet.
@@
  • We introduce GPUMemNet and GPUUtilNet, together with a fully reproducible estimation pipeline encompassing data generation scripts for MLP, CNN, and Transformer architectures, curated training datasets, and documented Jupyter notebooks, lowering the barrier for the community to extend and build upon our work.
@@
  • We provide a structured discussion of the key challenges, open limitations, and future directions in GPU memory and utilization estimation, with an emphasis on their implications for collocation and resource management in shared GPU clusters.
@@
The remainder of this paper is organized as follows: Section 2 reviews literature on GPU resource estimation and interference mitigation. Section 3 details the selected estimators, while Section 4 describes the dataset collection, problem formulation, feature engineering. Section 5 outlines the experimental setup, followed by a performance analysis in Section 6. Finally, Section 7 discusses limitations and future work, with concluding remarks in Section 8.
@@ 2. Related Work
GPU resource estimation and GPU resource management, specifically collocation management have received growing attention as shared GPU clusters suffer from underutilization (Jeon et al., 2019; Gao et al., 2024; Weng et al., 2022). We organize the related literature along two axes: static estimation approaches that predict memory and utilization prior to execution, and online, reactive approaches that manage interference dynamically during collocation.
@@ 2.1. GPU Memory and Utilization Estimation
Approaches to GPU memory estimation fall into three broad categories. Analytical methods derive memory estimates from model structure using closed-form accounting. Horus (Yeung et al., 2022) and DNNMem (Gao et al., 2020) are representative of this class; both require detailed architectural specifications and exhibit limited adaptability to transformer-based models and multi-GPU configurations, with DNNMem further lacking a publicly available implementation. LLMem (Kim et al., 2024) extends the analytical paradigm to distributed LLM fine-tuning, achieving low error rates, but remains tightly coupled to transformer decoder architectures, precluding generalization to other model families. Library-based CPU-side tools perform symbolic or formula-driven estimation without executing the model. xMem (Shi et al., 2025) utilizes the CPU to execute the initial iterations of a target task for a brief period to gather the necessary data for its estimation process. During this CPU execution, it collects detailed execution traces, including memory events and operator calls, allowing it to estimate peak memory requirements without needing a GPU. PyTorch FakeTensor (Contributors, 2025) propagates tensor shapes symbolically to estimate memory, while DeepSpeed (Microsoft DeepSpeed Team, 2023) provides built-in estimation utilities scoped to ZeRO-stage configurations; both require access to model source code or parameter counts and fail to account for framework-level overhead such as activation memory, gradient checkpointing, and caching allocator reservations. ML-based approaches attempt to overcome the rigidity of analytical methods; DNNPerf (Gao et al., 2023) employs graph neural networks to improve prediction accuracy, though it demands substantial computational resources and struggles to generalize to architectures unseen during training. Despite spanning three distinct paradigms, all methods share fundamental shortcomings: dependence on detailed model specifications or architecture internals, poor extensibility to unseen model families, and non-negligible inference latency that burdens the critical scheduling path. GPU utilization estimation remains comparatively understudied; Horus (Yeung et al., 2022) is among the few works to address it, yet its reliance on a small, curated model dataset limits its applicability to the diverse workloads encountered in production clusters.
@@ 2.2. Collocation and Interference Management
Online, reactive approaches have been proposed to manage interference that arises when multiple training jobs share a GPU. Gandiva (Xiao et al., 2018) supports collocation through time-slicing and introspective profiling of mini-batch progress rates, reverting packing decisions when significant slowdown is detected; however, this requires tight co-design between the ML framework and the resource manager, introducing substantial system intrusiveness. Gavel (Narayanan et al., 2020) adopts a round-based scheduling mechanism with a fixed round length to make heterogeneity-aware allocation decisions, but a static round granularity is inherently ill-suited for workloads with widely varying iteration times — too coarse for short jobs and too fine for long ones. Orion (Strati et al., 2024) achieves fine-grained, kernel-level collocation on a single GPU by classifying kernels as compute- or memory-bound through extensive offline profiling and intercepting kernel launches via an additional layer in the software stack; while effective, this approach incurs heavy profiling overhead and adds architectural complexity that hinders broader deployment. A common thread across these systems is that they treat interference reactively, after collocation has already been established, rather than proactively avoiding it through accurate pre-execution estimation. Furthermore, the dynamic nature of training workloads — where memory consumption and GPU utilization can shift substantially over the course of a job’s execution lifetime — remains largely unaddressed across both estimation and collocation literature, and we identify this as a critical frontier for future investigation.
3. Selected Estimators for This Study
3.1. GPU Memory Estimators
From the three classes of GPU memory estimators surveyed above, we select one available method per paradigm for empirical evaluation, the Horus analytical formula (Yeung et al., 2022), PyTorch FakeTensor (Contributors, 2025), and our own lightweight ML-based estimator, GPUMemNet. The reason for excluded estimators, for instance, DNNMem (Gao et al., 2020), which is not publicly available, LLMem (Kim et al., 2024), which is restricted to transformer decoder fine-tuning, and DeepSpeed’s estimator (Microsoft DeepSpeed Team, 2023), which is scoped exclusively to ZeRO-stage configurations and does not account for activation memory or framework overhead.
Analytical. The Horus formula (Yeung et al., 2022) estimates training memory via a lightweight analytical pass over model statistics, incurring negligible runtime overhead. However, as we demonstrate in Figure 1, it systematically overestimates memory consumption across MLP configurations due to its inability to account for framework-level optimizations such as activation reuse, dynamic memory management, and layer fusion. While this conservatism avoids OOM failures, it unnecessarily reduces collocation opportunities by over-reserving memory.
FakeTensor. PyTorch FakeTensor (Contributors, 2025) performs symbolic shape propagation, creating lightweight tensor surrogates that carry metadata — shape, dtype, device — without allocating real memory. This allows it to trace model execution and estimate memory without instrumentation of the training process itself. Evaluated across 2,030 model configurations from the TIMM (Wightman, 2019) library, FakeTensor consistently underestimates peak memory usage (Table 1), as it does not account for the caching allocator’s reserved memory, intermediate buffers, or optimizer states. A fixed safety margin (e.g., +4,GB) can partially compensate for this, providing a conservative bound suitable for collocation decisions. Its runtime overhead, while modest, is the most prominent among the three candidates and adds latency to the scheduling critical path. Additionally, FakeTensor fails on architectures that rely on unsupported operations, dynamic control flow, or CUDA-only kernels — including many Transformer variants, DLRM, and Mask R-CNN — limiting its practical coverage.
ML-based estimator GPUMemNet, our ML-based estimator, is trained on a curated synthetic dataset spanning MLP, CNN, and Transformer architecture families and formulates memory prediction as a classification task over discretized memory ranges. Unlike the analytical and library-based alternatives, it requires no model source code, no architectural specification beyond high-level configuration parameters, and introduces minimal inference latency. Its limitations and generalization behavior are analyzed in detail in Section 6.
P50 P80 P90 P95 P99 Max
Absolute error (GB) 1.16 1.97 2.82 3.84 6.82 (>> 8GB)
Estimation time (s) 0.94 1.49 1.84 2.28 3.47 4.97
3.2. GPU Utilization Estimator
Given that GPU utilization metrics remain largely understudied and the Horus dataset and models are currently unavailable, we propose a deep learning-based estimator for utilization metrics—specifically SMACT, SMOCC, and DRAMA—trained on the dataset we curated for GPU memory estimation.
4. Dataset and Problem Formulation
This section presents the curated dataset, covering data collection, processing, analysis, and the training and testing pipeline. We first outline the scripts and key considerations for constructing a valid dataset (Section 4.1), then justify the formulation of GPU memory and utilization estimation as a classification task (Section 4.2). We evaluated models of increasing complexity—from tree-based regressors to MLP- and Transformer-based architectures—using the latter to examine how layer sequences, activations, and parameters influence estimation performance. Finally, we describe the dataset and selected features for the MLP- and Transformer-based estimators (Section 4.3).
4.1. Data Collection, Processing, Analysis, and Training Pipeline
@@
To systematically collect GPU memory usage data, we developed a suite of scripts tailored to three different neural network architecture families: MLPs, CNNs, and Transformers. For each architecture, two main scripts are implemented:
  1. (1)
    Configurable Model Script:
@@
    This script defines a configurable neural network, supporting a flexible range of hyperparameters and architectural topologies (e.g., pyramid, hourglass, etc.). It ensures that only valid model configurations—feasible in terms of depth, width, and architecture—are generated. This validation guarantees the practicality of the network in real-world settings.
  2. (2)
    Launcher Script:
@@
    The launcher orchestrates data collection by uniformly sampling hyperparameters from predefined ranges. It launches a fixed number of configurations per run, with each configuration representing a unique data point. The parameters can be tweeked within the corresponding yaml file. Figure 2 shows inside the yaml file for MLP data collection. After configuring the parameters, running the script with python launcher_mlp.py, the data collection starts. The same goes for the two other as well.
@@
Each generated model is trained for a fixed, configurable duration of one minute, during which GPU and system resources are monitored via nvidia-smi(Corporation, [n. d.]), dcgmi(Corporation, 2024), and top (Documentation, [n. d.]). These tools capture GPU memory usage, utilization metrics (Yousefzadeh-Asl-Miandoab et al., 2023), power consumption, CPU load, and RAM usage. All raw monitoring logs and the model’s architectural summary are saved in dedicated per-configuration directories.

### v2

@@ Abstract.
Collocating deep learning training tasks improves GPU utilization but risks resource contention, severe slowdowns, and out-of-memory (OOM) failures. Accurate memory estimation is essential for robust collocation, and GPU utilization estimation — a key proxy for contention — enables interference-aware scheduling.
Existing GPU memory estimators span three paradigms – analytical models, CPU-side libraries, and ML-based estimators – each with distinct limitations: dependence on detailed model specifications, intrusive integration, poor generalization, and varying latency overhead. GPU heterogeneity further complicates estimation, as identical tasks can exhibit different memory footprints across hardware generations. GPU utilization remains comparatively understudied, further complicated by the non-additive utilization metrics and GPU heterogeneity.
We conduct a systematic analysis of representative estimators from each paradigm – Horus (Yeung et al., 2022), PyTorch FakeTensor (Contributors, 2025), and our lightweight ML-based estimator – evaluating accuracy, generalizability, and overhead. We construct a synthetic dataset spanning MLPs, CNNs, and Transformers with controlled architectural variations, and train MLP- and Transformer-based estimators for memory prediction, and experiment with utilization estimation. Our evaluation reveals key tradeoffs and validates estimators against real-world unseen models. Significant challenges remain: analytical models lack generalization and cannot easily be extended to new GPU architectures or accurately reflect memory optimization savings; CPU-side libraries impose intrusive integration overhead; and both analytical and ML-based estimators rely on model specifications or computation graphs, limiting generalization across diverse model architectures and GPU hardware variants. We release all datasets, tools, and artifacts to support further research.
@@ 1. Introduction
Accurate GPU memory estimation is critical for robust resource management in deep learning workloads, enabling appropriate GPU selection and safe task collocation without exceeding memory limits (Robroek et al., 2024). Estimating related metrics such as GPU utilization, as a proxy for interference, and memory bandwidth can further optimize collocation decisions and reduce performance degradations because of resource contention. However, estimation introduces latency into the critical path of resource manager decision-making. To mitigate this, estimation must be fast enough to meet scheduling deadlines, offloaded to machine learning frameworks (e.g., PyTorch) during model development, or pre-computed ahead of time while tasks are still queued and before they reach the scheduler.
@@
Deep learning training tasks are compute-intensive and demand substantial GPU memory based on model architecture, model size, input data size, and batch size, making efficient memory management essential. Maximizing GPU utilization is also necessary for energy efficiency (Jeon et al., 2019; Gao et al., 2024; Weng et al., 2022). Modern frameworks such as PyTorch and TensorFlow (Paszke et al., 2019; Abadi et al., 2016) employ optimization techniques like activation reuse and dynamic allocation, introducing unpredictability into memory usage patterns of training tasks. Consequently, simple estimation formulas based solely on numbers of parameters, gradients and activations often fail to produce accurate predictions. This complexity is further compounded by the fact that identical training tasks can exhibit markedly different memory footprints across different GPU hardware models (Albahar et al., 2022), owing to differences in software stacks, and hardware architecture. Beyond memory, GPU utilization estimation faces distinct challenges: hardware heterogeneity causes the same task to exhibit different utilization across different GPU models, and utilization metrics are non-additive, complicating collocation decisions.
@@
Existing GPU memory estimation approaches — spanning analytical methods (Yeung et al., 2022; Gao et al., 2020), library-based CPU-side tools (Contributors, 2025; Shi et al., 2025), and ML-based estimators (Gao et al., 2023) — suffer from over- or underestimation, dependence on detailed model specifications or computation graph, poor generalizability, or latency overhead on the critical scheduling/mapping path. GPU utilization estimation remains comparatively understudied, despite its use as a proxy for interference-aware collocation (Yeung et al., 2022) — a role complicated by its non-additive nature across collocated workloads. Reactive and/or profiling-based collocation systems such as Gandiva (Xiao et al., 2018), Gavel (Narayanan et al., 2020), and Orion (Strati et al., 2024) each address interference management under different assumptions, yet suffer from practical limitations: Gandiva requires intrusive ML framework modifications, Gavel incurs profiling overhead from its throughput estimator across accelerator types and its fixed profiling window may be too short to capture the behavior of long-running jobs or unnecessarily costly for short ones, and Orion requires costly offline per-kernel profiling.
In this paper, we analyze opportunities and limitations of the existing GPU memory and utilization estimators. Our contributions are as follows:
@@
  • We survey existing GPU memory estimation methods by categorizing them into three — analytical, library-based, and ML-based — and provide a comprehensive qualitative and quantitative comparison of representative techniques from each class: Horus (Yeung et al., 2022), PyTorch FakeTensor (Contributors, 2025), and our own GPUMemNet.
  • We introduce GPUMemNet and GPUUtilNet [^1] All artifacts, datasets, and reproducibility scripts are publicly available at https://github.com/itu-rad/GPUMemNet and https://github.com/itu-rad/GPUUtilNet., deep learning-based estimators for GPU memory consumption and utilization metrics, respectively. GPU memory estimators achieve up to 97% accuracy for MLPs and 82–87% for CNN and Transformer models, with limited generalization to real-world architectures such as ResNet, EfficientNet, and BERT. While GPUUtilNet reveals the difficulty of learning the patterns in the model architecture-level data.
@@
  • We release a fully reproducible estimation pipeline encompassing data generation scripts for MLP, CNN, and Transformer architectures, curated training datasets, and documented Jupyter notebooks, lowering the barrier for the community to extend and build upon our work.
@@
  • We provide a discussion of the key challenges, open limitations, and future directions in GPU memory and utilization estimation, with an emphasis on their implications for collocation and resource management in shared GPU clusters.
@@
The remainder of this paper is organized as follows: Section 2 reviews literature on GPU resource estimation and interference mitigation. Section 3 describes our ML-based GPU memory use and utilization estimators GPUMemNet and GPUUtilNet. Section 4 provides a performance analysis of a set of representative estimators. Finally, Section 5 discusses limitations and future work, with concluding remarks in Section 6.
@@ 2. Related Work
GPU resource estimation and GPU resource management, specifically collocation management have received growing attention as shared GPU clusters suffer from underutilization (Jeon et al., 2019; Gao et al., 2024; Weng et al., 2022). We organize the related literature along two axes: static estimation approaches that predict memory and utilization prior to execution (Section 2.1), and online, reactive approaches that manage interference dynamically during collocation (Section 2.2).
@@ 2.1. GPU Memory and Utilization Estimation
GPU memory use estimation methods fall into three categories.
Analytical methods derive memory estimates from model structure using closed-form accounting. Horus (Yeung et al., 2022) and DNNMem (Gao et al., 2020) are representative of this class; both require detailed architectural specifications and exhibit limited adaptability to transformer-based models and multi-GPU configurations, with DNNMem further lacking a publicly available implementation. LLMem (Kim et al., 2024) extends the analytical paradigm to distributed LLM fine-tuning, achieving low error rates, but remains tightly coupled to transformer decoder architectures, precluding generalization to other model families. A common limitation across analytical methods is their heavy reliance on detailed model specifications or computation graphs, which must either be explicitly provided by the user or obtained through intrusive coupling with the ML framework, and their limited extensibility — adapting to new configurations, such as multi-GPU settings, often requires revisiting and redesigning the underlying formulations rather than straightforward extension.
CPU-side Library-based tools perform symbolic or formula-driven estimation without requiring GPU execution. xMem (Shi et al., 2025) utilizes the CPU to execute the initial iterations of a target task for a brief period to gather the necessary data for its estimation process. During this CPU execution, it collects detailed execution traces, including memory events and operator calls, allowing it to estimate peak memory requirements without needing a GPU. However, xMem still requires partial CPU execution to collect memory traces, introducing overhead and raising representativeness concerns — in training workloads, peak memory may not manifest during the initial iterations due to optimizer state initialization, gradient accumulation, or warmup phases, potentially leading to underestimation. PyTorch FakeTensor (Contributors, 2025) propagates tensor shapes symbolically to estimate memory, requiring either model source code or an object instantiated from a checkpoint, while DeepSpeed (Microsoft DeepSpeed Team, 2023) provides built-in estimation utilities scoped exclusively to ZeRO-stage configurations, requiring either a model object or explicit parameter counts; both fail to account for framework-level overhead such as caching allocator reservations and dynamic memory allocation patterns that only emerge during actual execution.
Machine Learning (ML)-based approaches attempt to overcome the rigidity of analytical methods; DNNPerf (Gao et al., 2023) employs graph neural networks over the model’s computation graph, offering fine-grained operator-level accuracy, but at the cost of complex framework integration, inference latency that scales with graph complexity, and generalization challenges to unseen operator types or graph structures. More broadly, ML-based approaches are hindered by the difficulty of building sufficiently representative and diverse training datasets, making training, evaluation, and adaptation to new model families or hardware configurations a significant overhead.
Despite spanning three distinct paradigms, all methods introduce estimation overhead of varying magnitude, which can burden the critical path of resource manager decision-making.
In contrast, GPU utilization estimation remains comparatively understudied; Horus (Yeung et al., 2022) is among the few works to address it, yet its reliance on a small, curated model dataset limits its applicability to the diverse workloads encountered in production clusters.
@@ 2.2. Collocation and Interference Management
Online, reactive approaches have been proposed to manage interference that arises when multiple training jobs share a GPU. Gandiva (Xiao et al., 2018) supports collocation through time-slicing and introspective profiling of mini-batch progress rates, reverting packing decisions when significant slowdown is detected; however, this requires tight co-design between the ML framework and the resource manager, introducing intrusiveness to framework and resource management abstraction layers. Gavel (Narayanan et al., 2020) adopts a round-based scheduling mechanism with a fixed round length to make heterogeneity-aware allocation decisions, but a static round granularity is inherently ill-suited for workloads with widely varying iteration times — too coarse for short jobs and too fine for long ones. Orion (Strati et al., 2024) achieves fine-grained, kernel-level collocation on a single GPU by classifying kernels as compute- or memory-bound through extensive, and costly, offline profiling and intercepting kernel launches via an additional layer in the software stack. A common thread across these systems is that they treat interference reactively, after collocation has already been established, rather than proactively avoiding it through a priori estimation.
3. GPUMemNet & GPUUtilNet
To complement the estimation methods based on analytical formulas and libraries and strengthen the ML-based methods (described in Section 2.1), we first present a methodology to create a curated dataset and use it for ML-based estimators. This is necessary before we quantitatively evaluate the representative estimation methods from each category in Section 4. Section 3.1 first outlines the considerations and steps for constructing a valid dataset, then Section 3.2 justifies the formulation of GPU memory and utilization estimation as a classification task. Finally, Section 3.3 describes the dataset and selected features for the MLP- and Transformer-based estimators.
3.1. Steps for Data Curation
3.1.1. Considerations for Data Collection
To create a dataset with representative model features, we advocate for the following guidelines for data collection.
Architecture over model types. To ensure long-term relevance, we focus on architecture families (e.g., CNNs, Transformers) rather than specific models, which quickly become outdated. This means that the data collection scripts must create synthetic versions of the architecture families.
3.1.2. Scripts to Collect Data
@@
To systematically collect hardware utilization metrics while running the synthetic models created based on the considerations outlined above, we develop a suite of scripts tailored to three different neural network architecture families: MLPs, CNNs, and Transformers. For each architecture, two main scripts are implemented:
@@
Configurable Model Script defines a configurable neural network, supporting a flexible range of hyperparameters and architectural topologies (e.g., pyramid, hourglass, etc.). It ensures that only valid model configurations—feasible in terms of depth, width, and architecture—are generated. This validation guarantees the practicality of the network in real-world settings.
@@
Launcher Script orchestrates data collection by uniformly sampling hyperparameters from predefined ranges. It launches a fixed number of configurations per run, with each configuration representing a unique data point. The parameters can be tweaked within the corresponding yaml file. Figure 1 shows the yaml file for MLP data collection. After configuring the parameters, running the script with python launcher_mlp.py, the data collection starts.
@@
Each generated model is trained for a fixed, configurable duration of one minute, during which GPU and CPU resources are monitored via nvidia-smi(Corporation, [n. d.]), dcgmi(Corporation, 2024), and top (Documentation, [n. d.]). All raw monitoring logs and the model’s architectural summary are saved in dedicated per-configuration directories. Table 6, Table 5, Table 4, and Table 3 (All in Appendix) describe the curated dataset and its key parameters across model families.